An investment with an interest compounded continuously doubled itself in 12 years. What is the interest rate?
Answer by Theo(2978)

Formula for continuous compounding is:

FV = PA * e^(rt)

FV = future value
PA = present amount
e = scientific constant of 2.718281828.....
r = annual interest rate
t = time in years

In your equation,

FV = 2
PA = 1
e = 2.71828182818.....
r = r
t = 12

Equation becomes:

2 = 1 * e^(12r)

This becomes:

2 = e^(12r)

Take natural log of both sides (natural log is log to the base of e)

ln(2) = ln(e^(12r)

By laws of logarithms, this becomes:

ln(2) = 12r*ln(e)

ln(e) = 1

Equation becomes:

ln(2) = 12r

Divide both sides by 12 to get:

ln(2)/12 = r

Solve for r to get:

r = .057762265 * 100% = 5.7762265%

Using continuous compounding, an annual interest rate of 5.7762265% will double your money in 12 years.

Confirm by plugging into the original equation.

2 = 1 * e^(.057762265*12)

Simplify to get:

2 = 2 confirming the value of r is good.

You might try some of the LESSONS FROM ALGEBRA.COM if you have the time. There's a few in the financial area that might help you.